OpenStudy (anonymous):

someone tell me please im sure i did this question right but im told its wrong 13 1/2+8 1/4=?

OpenStudy (turingtest):

can you show me your work you I can point out your mistake?

OpenStudy (turingtest):

\[13\frac12+8\frac14=\frac{27}2+\frac{33}{4}=\frac{54}4+\frac{33}4=\frac{87}4\]do you see your mistake? do you follow this?

OpenStudy (anonymous):

can you simplify for me?

OpenStudy (anonymous):

yes i do sorry internet was acting up

OpenStudy (anonymous):

i do see what i did but not where i was as far off as my online class said i was i do follow that process

OpenStudy (turingtest):

well then screw your online class's way of doing it and do it however works best for you

OpenStudy (turingtest):

why dont you try to rewrite 87/4 as a mixed fraction? (I'm assuming taht is what you mean by "simplify"b ut that is not the right terminology)

OpenStudy (anonymous):

by simplify i mean to put it back into a mixed number with the integer and the fraction in its lowest common denominator

OpenStudy (turingtest):

I asked you to try it yourself first, if you don't mind ;) I wanna see where you made a mistake please show me your work, that will help me help you! 87/4 as a mixed fraction: what is it?

OpenStudy (anonymous):

had to restart hold on

OpenStudy (anonymous):

i got the same answer 21 3/4 right?

OpenStudy (anonymous):

heres my process 21x4 = 84 leaving 3 over 4 right?

OpenStudy (turingtest):

yeah that's right

OpenStudy (anonymous):

i knew it but what the heck? why does my class say the right answer is 40 something? i dont get that nvm its done it before thank you

OpenStudy (anonymous):

i really appreciate it

OpenStudy (turingtest):

teachers make mistakes books make mistakes humans make mistakes learn the material well, then just trust yourself that, in my opinion, is the best policy

OpenStudy (anonymous):

obviously im new to this could you add me im not sure how to do it

OpenStudy (anonymous):

thats a good oone true as can be too

OpenStudy (turingtest):

hover over my icon and click "become fan" then you can send me messages there are plenty here who can help you though

OpenStudy (anonymous):

thank you (i think) becoming a fan is friends here right?

OpenStudy (turingtest):

or write @TuringTest @b-van etc.

OpenStudy (anonymous):

in the post?

OpenStudy (turingtest):

yeah, that calls my attention who whatever problem you are working on dinner time, later bro!

OpenStudy (anonymous):

thanx again later!